OpenStudy (anonymous):

evaluate the limit of theta approaches zero of sin 3theta/ theta

OpenStudy (tkhunny):

REALLY BIG hint: \[\frac{\sin(3\cdot\theta)}{\theta} = \frac{3\cdot\sin(3\cdot\theta)}{3\cdot\theta} = 3\cdot\frac{\sin(3\cdot\theta)}{3\cdot\theta}\]